OpenStudy (anonymous):

If f(x) = 3a|4x – 4| – ax, where a is some constant, find f ′(1). a. 0 b. not enough information c. e d. 1 i don't understand this question because i thought the derivative doesn't exist at x=1 because there is a corner there.

OpenStudy (tkhunny):

It may or may not. Prove it. 1) Is it continuous? Does it have to be? 2) Are the left-and right- limits the same? Do they have to be?

OpenStudy (xapproachesinfinity):

first discuss lim f(x) at 1 then discuss continuity at 1

OpenStudy (xapproachesinfinity):

seems to me it is continuous at 1 since lim f =-a= f(1)

OpenStudy (xapproachesinfinity):

now you apply definition of derivative lim f(x)-f(1)/x-1 as x goes to 1 does it exist?

OpenStudy (anonymous):

i believe the limits are the same as x goes to 1 from the left and right side, so it does exist?

OpenStudy (xapproachesinfinity):

oh hold on i made a mistake you have to consider x>1 and x<1 x>1 ===> f(x)=3a(4x-4)-ax x<1 ===> f(x)=-3a(4x-4)-ax

OpenStudy (xapproachesinfinity):

now check the limits

OpenStudy (xapproachesinfinity):

an yes both limits are equal and they equal to f(1)=-a then f must be continuous for all x

OpenStudy (xapproachesinfinity):

now it the last part i said lim f(x)-f(1)/x-1 as x goes to 1 from both sides

OpenStudy (xapproachesinfinity):

if such limit exist then f must be differential at x=1

OpenStudy (xapproachesinfinity):

let's take x>1 \[\lim_{x\to 1^{+}}\frac{3a(4x-4)-ax+a}{x-1} = \lim 11a=11a\]

OpenStudy (xapproachesinfinity):

now you try x<1 \[\lim_{x\to 1^{-}}\]

OpenStudy (xapproachesinfinity):

i can already see that the f(1) doe not exist if you did that limit it won't be 11a

OpenStudy (xapproachesinfinity):

f'(1) i meant*

OpenStudy (anonymous):

i'm confused as to why you subtracted a from 3a(4x-4)-ax? did you use the equation f(x+h) +f(x)?

OpenStudy (anonymous):

oh i meant why you added not subtracted, sorry

OpenStudy (xapproachesinfinity):

yes i used the definition of derivative at 1 if f is differential at x=c, then the following limit must exist \[\lim_{x\to c}\frac{f(x)-f(c)}{x-c}\] this is the same as \[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

oh okay i see what you're doing

OpenStudy (xapproachesinfinity):

what does option C says? C.e do you mean Does not exists

OpenStudy (anonymous):

no i don't believe so. it just says e so i thought they were referring to Eulers number

OpenStudy (xapproachesinfinity):

oh ok because the f'(1) does not exist and your options don't have such thing i don't think i made a mistake

OpenStudy (anonymous):

yeah that was my whole thing. i thought it didn't exist so i thought i had done something wrong so i'm not sure what the question is asking

OpenStudy (xapproachesinfinity):

i gotta go i think we did it the correct way i don't know what the problem is

OpenStudy (anonymous):

okay no problem. thanks for the help!